the surface of a basketball is 576π cm². How far does this basketball travel if it makes 4 complete rolls?
Found 3 solutions by mananth, MathTherapy, ikleyn:
Answer by mananth(16946)


the surface of a basketball is 576π cm². How far does this basketball travel if it makes 4 complete rolls?
it will travel 4 times the perimeter around the diameter
area = 576pi = 4pi r^2
576/4 =r^2
144 =r^2
r = 12 ie the radius
diameter =12
Circumference = pi * d
12 pi
4 rolls
48 pi cm is the distance it travels

Answer by MathTherapy(10552)


the surface of a basketball is 576π cm². How far does this basketball travel if it makes 4 complete rolls?
His/her answer is WRONG, as usual! 


Answer by ikleyn(52781)

.

            Below you see my CORRECTED version of the solution by @mananth.

            Thanks to tutor @MathTherapy,  who detected and noticed an error in the solution by @mananth.



The surface of a basketball is 576π cm². How far does this basketball travel if it makes 4 complete rolls?

it will travel 4 times the circumference
area = 576pi = 4pi r^2
576/4 =r^2
144 =r^2
r = 12 the radius
diameter =24
Circumference = pi * d
24 pi
4 rolls
96 pi cm is the distance it travels.


/\/\/\/\/\/\/\/


You may ask me, for what reason I write and place here corrected versions of wrong solutions ?


            - I do it for future generations of students, who may look into this web-site,

            in order for they do not think that the current tutors all were incompetent . . .